# Measuring the Carbon Footprint of Cryptographic Privacy-Enhancing Technologies

### v1

@@ 3.3 Encrypted databases
Data. We run this experiment on five different websites: Wikipedia (Simplified English), New York Times, MDN, Mastodon Technical Blog, and xkcd. These websites are quite diverse: Wikipedia is an encyclopedia, New York Times a media, MDN a web developer documentation, Mastodon Blog is a blog, and xkcd is a minimal website publishing amusing comic strips. Wikipedia source files are publicly available, so we downloaded the “simplified English” archive from April 2007. We used the tool HTTrack to automatically download the static files (i.e., HTML, CSS, JS, and images) of the other websites.
Figure 7: Average energy consumption and carbon emissions of HTTP and HTTPS requests on five different websites (1000 requests).
Results. Figure 7 represents the average energy consumption and carbon footprint over 1000 HTTP(S) requests. The relative footprint increase ranges from 152% for NYTimes to 182% for xkcd.
@@
3.5 Encrypted emails
Protecting personal communications has long been a key objective for cryptographers. Over the decades, various solutions have emerged, ranging from encrypted emails to secure messaging apps like Signal. In this work, we focus on email encryption, as it is the most mature technology. Specifically, we examine the footprint of GPG encryption and signing, given that GPG is a widely used standard for securing email communication.
Software libraries. We use the GnuPG library (i.e., the reference open-source implementation for GPG). We benchmark 3 cipher suites: RSA, Elliptic Curve Cryptography (ECC), and ElGamal. Note that ElGamal only provides encryption so we combine it with DSA to sign messages. For each cipher suite, we used key sizes provided 128 bits of security [4].
Contrary to the other PETs, we do not have a non-private baseline to benchmark. Indeed, the non-private equivalent of email encryption is… no encryption (i.e., 0 cost). Thus, we present the encryption costs and put them into perspectives thanks to our previous experiments.
Data. We use the Enron email dataset to benchmark email encryption. We extract the 30109 emails contained in the sent mail folders of this dataset.
Results. Figure 8 presents the results. First, the encryption is systematically slightly more expensive than the signature for RSA and ECC. Second, RSA is slightly less expensive than ECC. Finally, ElGamal encryption is nearly ten times more expensive than RSA and ECC.
For ElGamal, remember that we use DSA for signatures. The figure shows that DSA is about twice as expensive as RSA and ECC signatures.
Figure 8: Average energy consumption and carbon emissions of GPG email encryption and signing (30K emails).
To put these results into perspectives, we can compare them to secure ML inference. Based on our results, the footprint of an email encryption is then $1000\times$ smaller than the footprint of an encrypted random forest inference.
Discussions. Our comparison shows comparable computational costs for RSA and ECC. However, ECC has another advantage over RSA: smaller key sizes. ECC could reduce (by a couple of kilobytes) the communication costs related to key and signature exchanges. As explained in Section 2.3, there is no consensus (yet) on the overall footprint of network communications, so we cannot include this factor in our comparison.
Finally, email encryption is becoming less and less popular; in favor of secure messaging apps that are more accessible. Such secure messaging systems rely heavily on symmetric-key cryptography, which is known to be more efficient than public-key cryptography. Secure messaging apps could then provide even smaller footprints than those of email encryption.
However, such messaging systems also require expensive protocols to set up communications. For example, half the operation costs of Signal Messenger is caused by the registration process [43]. An LCA of Signal should then include these expensive operations and not only the cheap symmetric-key encryption.
Our analysis of email encryption provides valuable baseline, but the analysis of complex messaging systems such as Signal represents a promising future work. Such analysis is particularly challenging as some protocols used by Signal require cryptographic hardware like secure enclaves whose environmental footprint is harder to estimate (as discussed in Section 4.1).
@@ 4 Orthogonal discussions
4.1 The hidden cost of cryptographic hardware
@@
However, the LCA must include the manufacturing impact, as this specialized hardware adds a hardware overhead in our taxonomy. Since manufacturing accounts for most of the ICT sector’s carbon footprint [21], using cryptographic accelerators does not necessarily reduce the overall carbon footprint.
Trusted Execution Environments (TEEs) [39] are another example of cryptographic hardware with an environmental cost that should not be ignored. TEEs are now widely used in efficient secure protocols because they allow operations on confidential data without relying on expensive primitives like homomorphic encryption. For example, Signal uses TEEs in its Contact Discovery protocol.
@@
4.2 Privacy-Carbon-Functionality trilemma
Our work initially examined the relationship between privacy and carbon emissions in an isolated manner. However, our experiments indirectly revealed a broader trilemma involving privacy, carbon emissions, and functionality.
The carbon footprint of cryptographic PETs is tied to their functionality; more complex functionalities generally require more computationally expensive cryptographic primitives, leading to higher emissions. This phenomenon is evident in Figure 2. While encryption always increases the footprint, the extent of this increase varies significantly across ML models. For instance, logistic regression exhibits a 100-fold increase, whereas encrypted neural networks result in a 100,000-fold increase. Consequently, simpler models like logistic regression lead to a smaller footprint.
@@ 4.2 Privacy-Carbon-Functionality trilemma
4.3 Mitigating carbon footprint overhead via decentralization
Besides functionality, trust is another powerful leverage to optimize the privacy-carbon trade-off. The PET literature usually considers two edge cases: (1) a party trusted by all users requiring no PET and (2) a zero-trust world requiring (potentially) expensive PETs.
@@ 4.3 Mitigating carbon footprint overhead via decentralization
[…]
Decentralization “à la Mastodon” is then a promising direction to enhance privacy while avoiding significant carbon footprint overheads.
@@ 5 Conclusion
Our work studied the environmental footprint of cryptographic PETs. We first introduced a standardized methodology for measuring the carbon footprint of PETs; demonstrating it on five cryptographic PETs. Our results highlight (PET-induced) carbon footprint increases ranging from $2\times$ to $100,000\times$. Our findings provide essential data to assist decision-makers assessing the privacy-carbon trade-offs inherent in different ICTs.
Runtime has always been a natural metric in PET papers, and recent works [37] also mention monetary costs on Amazon Web Services. Even though such costs are relatable to practitioners, they could be completed with energy consumption (and possibly carbon emission) measurements. While monetary costs are valuable to economical actors, the environmental footprint is highly informative for the society as a whole (e.g., journalists or policymakers).
Future directions
Our work marks a first step in the carbon footprint analysis of PETs, but significant work is still necessary to analyze the footprint of all existing PETs (especially complex systems such as Signal Messenger or Tor).

### v2

@@ 3.4 TLS
Data. We run this experiment on five different websites: Wikipedia (Simplified English), New York Times, MDN, Mastodon Technical Blog, and xkcd. These websites are quite diverse: Wikipedia is an encyclopedia, New York Times a media, MDN a web developer documentation, Mastodon Blog is a blog, and xkcd is a minimal website publishing amusing comic strips. Wikipedia source files are publicly available, so we downloaded the “simplified English” archive from April 2007. We use the tool HTTrack to automatically download the static files (i.e., HTML, CSS, JS, and images) of the other websites.
Figure 5: Average energy consumption of HTTP and HTTPS requests on five different websites (1000 requests).
Results. Figure 5 represents the average energy consumption over 1000 requests. The relative increase ranges from 152% for NYTimes to 182% for xkcd.
@@ 4 Orthogonal discussions
4.1 The hidden footprint of cryptographic hardware
@@
However, the LCA must include the manufacturing impact, as this specialized hardware requires the production of additional hardware (in addition to the standard CPU). Since manufacturing accounts for most of the ICT sector’s carbon footprint [17], using cryptographic accelerators does not necessarily reduce the overall carbon footprint.
Trusted Execution Environments (TEEs) [33] are another example of cryptographic hardware with an indirect environmental footprint that should not be ignored. TEEs are now widely used in efficient secure protocols because they allow operations on confidential data without relying on expensive primitives like homomorphic encryption. For example, Signal uses TEEs in its Contact Discovery protocol.
@@
4.2 Privacy-Energy-Functionality trilemma
Our work initially examined the relationship between privacy and energy emissions in an isolated manner. However, our experiments indirectly revealed a broader trilemma involving privacy, energy emissions, and functionality.
The energy consumption of cryptographic PETs is tied to their functionality; more complex functionalities generally require more computationally expensive cryptographic primitives, leading to higher emissions. This phenomenon is evident in Figure 1. While encryption always increases the consumption, the extent of this increase varies significantly across ML models. For instance, logistic regression exhibits a 100-fold increase, whereas encrypted neural networks result in a 100,000-fold increase. Consequently, simpler models like logistic regression lead to a smaller consumption.
@@ 4.2 Privacy-Energy-Functionality trilemma
4.3 Mitigating overhead via decentralization
Besides functionality, trust is another powerful leverage to optimize the privacy-energy trade-off. The PET literature usually considers two edge cases: (1) a party trusted by all users requiring no PET and (2) a zero-trust world requiring (potentially) expensive PETs.
@@ 4.3 Mitigating overhead via decentralization
We can formulate the Mastodon threat model as follows: each user trusts one specific server (among all possible servers), and they do not trust anyone else. Such decentralized trust avoids expensive cryptographic operations because each server can perform plaintext operations on the personal data of its users. For instance, on Peertube (a “decentralized Youtube”), each server can offer ML-based video recommendations to its users without relying on encrypted ML.
Decentralization “à la Mastodon” is then a promising direction to enhance privacy while avoiding significant energy consumption overheads.
@@
However, one may argue that decentralized protocols require infrastructure redundancy as it requires many servers. For example, Mastodon is a decentralized social media composed of 8K servers (in September 2025). But, it does not imply that Mastodon requires 10K times more servers than a centralized social media. Indeed, centralized social media commonly rely on a global network of servers. Thus, both decentralized and centralized social media rely on a network of servers; the only difference is in the server ownership. Thus, the subtle question of infrastructure and communication overheads in decentralized protocols must be careful studied because even large-scale centralized services rely on distributed systems creating redundancies and overheads. This calls for dedicated future work (e.g., comparing Matrix and Signal protocols in secure messaging).
@@ 5 Conclusion
Our work studied the energy consumption of three cryptographic PETs with different levels of maturity. Our results highlight PET-induced energy consumption increases ranging from $2\times$ (for TLS) to $100{,}000\times$ (for FHE). This wide range also suggests a relationship between maturity and efficiency: the most mature technology is the most energy-efficient.
However, this correlation should not hide the fact that FHE remains several orders of magnitude more expensive than its plaintext equivalent. While performance optimization is a central goal in cryptography research, this raises a key question: what is the optimal performance? Modern FHE cryptosystems rely on inherently expensive primitives, unlike TLS, which makes reaching comparable efficiency difficult, if not impossible. This observation should not discourage research on FHE, but it illustrates a potential broader debate for the cryptography community about the trade-off between privacy and environmental footprint.